OpenStudy (anonymous):

Multiply (1.2 ⋅ 1028) ⋅ (3 ⋅ 10−19). Express the answer in scientific notation.

OpenStudy (anonymous):

help please

OpenStudy (phi):

do you mean \[ (1.2 ⋅ 10^{28}) ⋅ (3 ⋅ 10^{−19}). \]? when multiplying, you can reorder the numbers. For example \[ 1.2 \cdot 3 \cdot (10^{28} \cdot 10^{−19}) \] I assume you can figure out 1.2*3 ? for the 10 to the exponent, the product is the *sum of the exponents" to the base 10